# SPARKE: Scalable Prompt-Aware Diversity Guidance in Diffusion Models via RKE Score

### v1

@@ 1 Introduction
Diffusion models [1, 2, 3] have rapidly become a prominent class of generative models, achieving state-of-the-art results in several generative modeling tasks [2, 4, 5]. Notably, their ability to produce intricate and realistic image and video data has significantly advanced various content creation pipelines [6, 7, 8, 9]. Despite these successes, ensuring sufficient diversity in generated outputs remains an area of active research, particularly in prompt-guided diffusion models, where the diversity of generated samples needs to be evaluated and optimized while considering the variability of the input prompts.
@@
Diversity in prompt-guided sample generation also plays a critical role in ensuring fairness and lack of mode collapse in response to a certain prompt category, implying that the outputs do not collapse in response to a group of prompts with similar components. Therefore, an adequately diverse prompt-guided generation model is supposed to provide varied and balanced outputs conditioned on each prompt category. For example, in geographically diverse image synthesis, the generation model is supposed to represent a broad spectrum of regional styles, preventing overrepresentation of dominant patterns while ensuring fair coverage across different locations [10]. Similarly, in data augmentation for visual recognition, generating a semantically varied set of samples enhances the model’s generalizability and reduces biases in prompt-guided sample creation [11, 12].
Figure 1: Overview of the proposed SPARKE method in generating images at different iterations in comparison to the vanilla Stable Diffusion-XL (SDXL) [5] model. We also compare the conditional-RKE guidance in SPARKE with the baseline Vendi Score guidance [10] (unconditional, in latent space).
@@
A recently explored strategy for diversity-aware generation in diffusion models is to utilize a diversity score for guiding the sample generation. One such method is the contextualized Vendi score guidance (c-VSG) [10]. The c-VSG approach builds upon the Vendi diversity measure [13], which is a kernel matrix-based entropy score evaluating diversity in generative models. The numerical results in [10] demonstrate the effectiveness of regularizing the Vendi score in guiding diffusion models toward more diverse sample distributions over multiple rounds of data generation. Another method, CADS [14] tends to improve sample diversity through the perturbation of conditional inputs, typically resulting in an increased Vendi score.
@@
Although the mentioned diversity-score-based diffusion sampling methods apply smoothly to the unconditional sample generation without a varying input prompt, a scalable prompt-aware extension to prompt-guided diffusion models remains a challenge. As discussed in [10], one feasible framework extension to prompt-guided generation can leverage a pre-clustering of prompts to apply the Vendi diversity guidance separately within every prompt cluster. However, in successive queries to prompt-guided generative models, prompts often vary in subtle details, making their pre-clustering into a fixed number of hard clusters challenging. In addition, relying on small clusters of highly similar prompts could lead to limited diversity across prompt clusters, due to missing the partial similarities of the prompts in different clusters.
@@
In this work, we propose Scalable Prompt-Aware Rény Kernel Entropy Diversity Guidance (SPARKE) in diffusion models, applying the conditional entropy-based diversity score family [15] of the latent representation in the latent diffusion models (LDMs). As illustrated in Figure 1, SPARKE directly incorporates prompts into the diversity calculation without pre-clustering. This extension enables a more dynamic control over the diversity calculation with varying prompts, which can be interpreted as an adaptive kernel-based similarity evaluation without grouping the prompts into clusters. According to the conditional entropy guidance in SPARKE, we update the kernel matrix of generated data by taking the Hadamard product with the kernel matrix of input prompts. Applying an appropriate kernel function for text prompts, the updated similarity matrix will assign a higher weight to the pairwise interaction of samples with higher prompt similarity.
@@
Figure 2: Qualitative Comparison of samples generated by base PixArt-$\Sigma$ and Stable Diffusion XL vs. SPARKE diversity guided sampling.
To improve scalability while preserving diversity-aware generation, we propose applying the order-2 Renyi matrix-based entropy, defined as the RKE Score [16], in the SPARKE approach. This RKE-based formulation replaces eigenvalue decomposition with a Frobenius norm-based entropy measure, which significantly reduces computational overhead. As a result, the complexity of entropy estimation reduces to $\mathcal{O}(n^{2})$, while the gradient-based optimization in guided diffusion models can be performed even more efficiently in $\mathcal{O}(n)$ time due to the canceled terms with zero gradients. By integrating these computational improvements, we introduce SPARKE framework, which maintains the advantages of prompt-conditioned diversity guidance while making large-scale sample generation feasible.
@@
  • We propose the application of entropy guidance in the latent space of latent diffusion models (LDMs), improving the efficiency and performance in the entropy-based diversity guidance.
  • We test our method on several state-of-the-art diffusion models (see Figures 2 and 8) and text-to-image benchmarks, indicating its ability to improve sample diversity and computational efficiency.
@@ 2 Related Works
Standard and Latent Diffusion Models. Diffusion generative models [1, 2, 17, 3] learns to reverse an iterative noising process, effectively estimating the gradient of the data log-density (the Stein score [18]) to generate new data samples. This approach has shown remarkable capabilities in synthesizing high-fidelity images [6, 19, 20, 7]. Despite the impressive results, a primary limitation of diffusion models was the substantial computational cost, particularly when operating directly in high-dimensional spaces like pixel space. To address this challenge, Latent Diffusion Models (LDMs) [4, 21] perform the forward and denoising processes within an encoded latent space $\bm{z}=\mathcal{E}(\bm{x})$, enabling high-quality images such as Stable Diffusion [4, 5] and video [22, 23, 8] synthesis at an unprecedented scale.
@@
Conditional Generation with Guidance. The ability to control generative processes with specific conditions is increasingly crucial for practical applications, based on conditional inputs like text-guided [20, 24, 25], class labels [6], style images [26, 27], or human motions [28], etc. Methods for conditional generation with guidance are categorized as either training-based or training-free. Training-based approaches either learn a time-dependent classifier that guides the noisy sample $\bm{x}_{t}$ towards the condition $\bm{y}$ [6, 24, 29, 25], or directly train the conditional denoising model $\bm{\epsilon}_{\theta}(\bm{x}_{t},t,\bm{y})$ via few-shot adaptation [26, 4, 30]. In contrast, training-free guidance aims at zero-shot conditional generation by leveraging a pre-trained differentiable target predictor without requiring any training. This predictor can be a classifier, loss function, or energy function quantifying the alignment of a generated sample with the target condition [31, 32, 33, 34]. Our work can be included as a training-free guidance approach that applies conditional entropy scores guidance to enhance the diversity of generated samples.
Diversity Evaluation of Generative Models. Evaluation metrics for generative models are broadly categorized into reference-based and reference-free approaches. Reference-based metrics, such as Precision/Recall [35, 36], and Density/Coverage [37] measure quality and diversity by comparing generated samples to real data. Novelty assessment has also been explored via metrics like the Rarity Score [38], KEN [39], and FINC [40, 41]. In contrast, reference-free metrics evaluate generations without relying on real data [16, 13, 42, 43]. [15, 44] proposed reference-free metrics for conditional generative models. In particular, [15] introduced a family of scores called Conditional-Vendi, which we use in this work for prompt-aware diversity guidance.
@@
Guidance for Improving Diversity. A common strategy in diffusion-based generative modeling is the use of guidance mechanisms to balance quality and diversity [45, 46]. For example, classifier-free guidance methods [46] considerably enhance prompt alignment and image quality but often compromise diversity due to overly deterministic conditioning. Several recent works have attempted to address this diversity challenge. [47] introduced a diversity-promoting strategy by explicitly sampling from low-density regions of the data manifold, encouraging diversity, however, their method operates directly in pixel space, posing challenges in adapting it effectively to latent diffusion frameworks. Another line of work is fine-tuning. In [48], the authors provide a finetuning method using Reinforcement Learning to improve the diversity of generated samples using an image-set-based diversity reward function.
@@
Recent works tackle this problem in the denoising phase. The CADS framework [14] shows that adding Gaussian noise to the conditioning signals during inference increases sample diversity. Their annealing schedule provides a balance between diversity and quality. Particle Guidance (PG) [49] employs non-IID sampling from the joint distribution defined by a diffusion model combined with a potential function that maximizes pairwise dissimilarity across all samples, independent of semantic context. Similar to PG, ProCreate [50] utilizes log energy calculated from embeddings provided by the DreamSim network to identify highly similar images, and uses maximum embedding-space distances across all generated objects. A concurrent work, SPELL [51], adds repellency terms during sampling, preventing the predicted samples from landing too close within the batch. These methods maximize across all samples without considering the prompts. On the other hand, our method focuses on a prompt-aware approach, dynamically conditioning diversity guidance on input prompts.
@@
The recent work [10] introduces contextualized Vendi Score Guidance (c-VSG), enhancing generative diversity during the denoising process using Vendi Score [13, 42]. Their approach, however, requires the same prompts, restricting its applicability across diverse prompt scenarios. On the other hand, our method leverages the RKE Score [16] to significantly enhance computational efficiency and sample complexity. Additionally, we propose prompt-aware guidance inspired by [15], enabling adaptive soft-clustering and effective conditioning on semantically distinct prompts. Moreover, unlike [10], which applies guidance based on latent encoded features of reference images, our strategy directly applies guidance within the diffusion model’s latent space, saving computational costs and enabling a prompt-aware diversity improvement.
@@ 3 Preliminaries
3.1 Kernel Matrices and RKE Diversity Scores
@@
The Rényi Kernel Entropy (RKE) diversity score is defined and analyzed by  [16] as the inverse Frobenius norm-squared of the trace-normalized kernel matrix:
$\mathrm{RKE}\bigl{(}x_{1},\ldots,x_{n}\bigr{)}:=\Bigl{\|}\frac{K_{X}}{\mathrm{Tr}(K_{X})}\Bigr{\|}^{-2}_{F},$ (2)
where $\mathrm{Tr}(\cdot)$ is the trace operator and $\|\cdot\|_{F}$ denotes the Frobenius norm. As theoretically shown in [16], the above RKE score can be regarded as a mode count of a mixture distribution with multiple modes.
To extend this entropy-based diversity scores to conditional (i.e., prompt-aware) diversity measurement for prompt-guided generative models, [15] propose the application of conditional kernel matrix entropy measures, resulting in the following definition of order-$2$ Conditional-RKE given the Hadamard product $\frac{1}{\mathrm{Tr}(K_{X}\odot K_{T})}K_{X}\odot K_{T}$ of the output data $K_{X}$ kernel matrix and input prompt kernel $K_{T}$ matrix:
$\mathrm{Conditional}\text{-}\mathrm{RKE}\bigl{(}x_{1},\ldots,x_{n};t_{1},\ldots,t_{n}\bigr{)}\,:=\,\Bigl{\|}\frac{{K}_{T}}{\mathrm{Tr}({K}_{T})}\Bigr{\|}^{2}_{F}\cdot\,\Bigl{\|}\frac{{K}_{T}\odot{K}_{X}}{\mathrm{Tr}({K}_{T}\odot{K}_{X})}\Bigr{\|}^{-2}_{F}$ (3)
Note that here the Hadamard product of the prompt $(t_{1},\ldots t_{n})$ kernel matrix $K_{T}$ and output $(x_{1},\ldots x_{n})$ kernel matrix $K_{X}$ is defined as the elementwise product of the matrices, which is guaranteed to be a PSD matrix by the Schur product theorem,
${K}_{T}\odot K_{X}\,:=\,\begin{bmatrix}k(t_{1},t_{1})\cdot k(x_{1},x_{1})&\cdots&k(t_{1},t_{n})\cdot k(x_{1},x_{n})\vspace{2mm}\\ \vdots&\ddots&\vdots\vspace{2mm}\\ k(t_{n},t_{1})\cdot k(x_{n},x_{1})&\cdots&k(t_{n},t_{n})\cdot k(x_{n},x_{n})\end{bmatrix}$ (4)
@@ 3.2 Conditional Latent Diffusion Models
Latent Diffusion Models (LDMs) [4] can achieve scalability and efficiency by performing the diffusion process within a compressed latent space, rather than directly in pixel space. A pre-trained autoencoder is used in LDMs, consisting of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. The encoder $\mathcal{E}$ maps high-dimensional images $\bm{x}\in\mathbb{R}^{H\times W\times C}$ to a lower-dimensional latent representation $\bm{z}_{0}=\mathcal{E}(\bm{x})\in\mathbb{R}^{h\times w\times c}$, while the decoder $\mathcal{D}$ reconstructs the images $\tilde{\bm{x}}=\mathcal{D}(\bm{z}_{0})$ from a denoised latent variable $\bm{z}_{0}$. The forward diffusion process applies in the latent space as $\bm{z}_{t}=\sqrt{\alpha_{t}}\bm{z}_{0}+\sqrt{1-\alpha_{t}}\bm{\epsilon}_{t}$, where $\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathit{\bm{I}})$ and $\alpha_{t}\in$ are predefined noise schedule parameters. Then the conditional LDMs with conditions $\bm{y}$ are trained to predict the noise $\bm{\epsilon}_{\theta}(\bm{z}_{t},t,\bm{y})$ at each time step $t$, also learn the score of $p_{t}(\bm{z_{t}}|\bm{y})$ [17, 3]:
$\min_{\theta}\mathbb{E}_{\bm{z}_{t},\bm{\epsilon}_{t},t,\bm{y}}\left[\left\|\bm{\epsilon}_{\theta}(\bm{z}_{t},t,\bm{y})-\bm{\epsilon}_{t}\right\|_{2}^{2}\right]=\min_{\theta}\mathbb{E}_{\bm{z}_{t},\bm{\epsilon}_{t},t,\bm{y}}\left[\left\|\bm{\epsilon}_{\theta}(\bm{z}_{t},t,\bm{y})+\sqrt{1-\alpha_{t}}\nabla_{\bm{z}_{t}}\log p_{t}(\bm{z}_{t}|\bm{y})\right\|_{2}^{2}\right].$ (5)

### v2

# SPARKE: Scalable Prompt-Aware Diversity and Novelty Guidance in Diffusion Models via RKE Score

@@ 1 Introduction
Diffusion models sohl2015deep ; ho2020denoising ; song2020score have rapidly become a prominent class of generative models, achieving state-of-the-art results in several generative modeling tasks ho2020denoising ; rombach2022high ; podell2024sdxl . Notably, their ability to produce intricate and realistic image and video data has significantly advanced various content creation pipelines dhariwal2021diffusion ; ho2022cascaded ; blattmann2023stable ; luo2021diffusion . Despite these successes, ensuring sufficient diversity in generated outputs remains an area of active research, particularly in prompt-guided diffusion models, where the diversity of generated samples needs to be evaluated and optimized while considering the variability of the input prompts.
@@
Diversity in prompt-guided sample generation also plays a critical role in ensuring fairness and lack of mode collapse in response to a certain prompt category, implying that the outputs do not collapse in response to a group of prompts with similar components. Therefore, an adequately diverse prompt-guided generation model is supposed to provide varied and balanced outputs conditioned on each prompt category. For example, in geographically diverse image synthesis, the generation model is supposed to represent a broad spectrum of regional styles, preventing overrepresentation of dominant patterns while ensuring fair coverage across different locations askari2024improving . Similarly, in data augmentation for visual recognition, generating a semantically varied set of samples enhances the model’s generalizability and reduces biases in prompt-guided sample creation miao2024training ; shorten2021survey .
@@
A recently explored strategy for diversity-aware generation in diffusion models is to utilize a diversity score for guiding the sample generation. One such method is the contextualized Vendi score guidance (c-VSG) askari2024improving . The c-VSG approach builds upon the Vendi diversity measure friedman2023vendi , which is a kernel matrix-based entropy score evaluating diversity in generative models. The numerical results in askari2024improving demonstrate the effectiveness of regularizing the Vendi score in guiding diffusion models toward more diverse sample distributions over multiple rounds of data generation. Another method, CADS sadat2024cads tends to improve sample diversity through the perturbation of conditional inputs, typically resulting in an increased Vendi score.
@@
Although the mentioned diversity-score-based diffusion sampling methods apply smoothly to the unconditional sample generation without a varying input prompt, a scalable prompt-aware extension to prompt-guided diffusion models remains a challenge. As discussed in askari2024improving , one feasible framework extension to prompt-guided generation can leverage a pre-clustering of prompts to apply the Vendi diversity guidance separately within every prompt cluster. However, in successive queries to prompt-guided generative models, prompts often vary in subtle details, making their pre-clustering into a fixed number of hard clusters challenging. In addition, relying on small clusters of highly similar prompts could lead to limited diversity across prompt clusters, due to missing the partial similarities of the prompts in different clusters.
Figure 1: Overview of the proposed SPARKE method in generating images at different iterations in comparison to the vanilla Stable Diffusion-XL podell2024sdxl model. We also compare the conditional-RKE guidance in SPARKE with the baseline Vendi Score guidance askari2024improving (unconditional, in latent space).
@@
In this work, we propose Scalable Prompt-Aware Rény Kernel Entropy Diversity Guidance (SPARKE) in diffusion models, applying the conditional entropy-based diversity score family jalali2024conditional of the latent representation in the latent diffusion models (LDMs). As illustrated in Figure 1, SPARKE directly incorporates prompts into the diversity calculation without pre-clustering. This extension enables a more dynamic control over the diversity calculation with varying prompts, which can be interpreted as an adaptive kernel-based similarity evaluation without grouping the prompts into clusters. According to the conditional entropy guidance in SPARKE, we update the kernel matrix of generated data by taking the Hadamard product with the kernel matrix of input prompts. Applying an appropriate kernel function for text prompts, the updated similarity matrix will assign a higher weight to the pairwise interaction of samples with higher prompt similarity.
@@
Figure 2: Comparison of latent entropy-based diversity guidance (ours) vs. ambient entropy diversity guidance in Latent Diffusion Models (LDMs). The experiment is performed with the SD-XL LDM.
Figure 3: Qualitative Comparison of samples generated by the base latent diffusion model (LDM), PixArt-$\Sigma$, and Stable Diffusion XL, vs. LDM guided via our proposed SPARKE guidance.
To improve scalability while preserving diversity-aware generation, we propose applying the order-2 Renyi matrix-based entropy, defined as the RKE Score jalali2023information , in the SPARKE approach. This RKE-based formulation replaces eigenvalue decomposition with a Frobenius norm-based entropy measure, which significantly reduces computational overhead. As a result, the complexity of entropy estimation reduces to $\mathcal{O}(n^{2})$, while the gradient-based optimization in guided diffusion models can be performed even more efficiently in $\mathcal{O}(n)$ time due to the canceled terms with zero gradients. By integrating these computational improvements, we introduce SPARKE framework, which maintains the advantages of prompt-conditioned diversity guidance while making large-scale sample generation feasible.
@@
  • We propose the application of entropy guidance in the latent space of latent diffusion models, improving the efficiency and performance in the entropy-based diversity guidance (See Figure 2).
  • We test our method on several state-of-the-art diffusion models (see Figures 3) and text-to-image benchmarks, indicating its ability to improve sample diversity and computational efficiency.
@@ 2 Related Works
Standard and Latent Diffusion Models. Diffusion generative models sohl2015deep ; ho2020denoising ; song2019generative ; song2020score learns to reverse an iterative noising process, effectively estimating the gradient of the data log-density (the Stein score stein1972bound ) to generate new data samples. This approach has shown remarkable capabilities in synthesizing high-fidelity images dhariwal2021diffusion ; nichol2021improved ; kim2022diffusionclip ; ho2022cascaded . Despite the impressive results, a primary limitation of diffusion models was the substantial computational cost, particularly when operating directly in high-dimensional spaces like pixel space. To address this challenge, Latent Diffusion Models (LDMs) rombach2022high ; vahdat2021score perform the forward and denoising processes in an encoded latent space, enabling high-quality images such as Stable Diffusion rombach2022high ; podell2024sdxl and video he2022latent ; blattmann2023align ; blattmann2023stable synthesis at a large scale.
@@
Conditional Generation with Guidance. The ability to control generative processes with specific conditions is increasingly crucial for practical applications, based on conditional inputs like text-guided kim2022diffusionclip ; nichol2021glide ; liu2023more , class labels dhariwal2021diffusion , style images mou2024t2i ; zhang2023adding , or human motions tevet2022human , etc. Methods for conditional generation with guidance are categorized as either training-based or training-free. Training-based approaches either learn a time-dependent classifier that guides the noisy sample $\boldsymbol{x}_{t}$ towards the condition $\boldsymbol{y}$ dhariwal2021diffusion ; nichol2021glide ; zhao2022egsde ; liu2023more , or directly train the conditional denoising model $\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t,\boldsymbol{y})$ via few-shot adaptation mou2024t2i ; rombach2022high ; ruiz2023dreambooth . In contrast, training-free guidance aims at zero-shot conditional generation by leveraging a pre-trained differentiable target predictor without requiring any training. This predictor can be a classifier, loss function, or energy function quantifying the alignment of a sample with the target condition he2023manifold ; bansal2023universal ; yu2023freedom ; ye2024tfg . Our work can be included as a training-free guidance approach that applies conditional entropy scores guidance to enhance the diversity of samples.
Quantifying Diversity and Novelty. Diversity is quantified using both reference-based kynkanniemi2019 ; naeem2020reliable and reference-free metrics. Reference-free metrics include the Vendi Score friedman2023vendi ; ospanov2024towards ; ospanov2024statvendi , the RKE score jalali2023information for unconditional models, and the Conditional-RKE jalali2024conditional and Scendi Ospanov_2025_ICCV scores for conditional models. Also, the diversity metrics have been extended to online and distributed model selection tasks rezaei2025mixtureucb ; hu2025onlineeval ; hu2025pakucb ; wang2023distributedKID . For novelty, prior work jiralerspong2023feature ; zhang2024interpretable ; zhang2025finc analyzes how generated samples differ from a reference model, with zhang2024interpretable ; zhang2025finc proposing a spectral method to measure the entropy of novel modes. jalali2025spec ; gong2025kernelbasedunsupervisedembeddingalignment ; gong2025boosting ; wu2025when also introduce kernel-based methods to compare and align two embeddings. In this work, we propose a novelty guidance approach that operates with respect to a reference dataset.
@@
Guidance for Improving Diversity. A common strategy in diffusion-based generative modeling is the use of guidance mechanisms to balance quality and diversity sadat2025no ; ho2022classifier . For example, classifier-free guidance methods ho2022classifier considerably enhance prompt alignment and image quality but often compromise diversity due to overly deterministic conditioning. Several works have attempted to address this diversity challenge. To encourage diversity, Sehwag_low_density introduced a strategy that samples from the data manifold’s low-density regions, however, their method operates directly in pixel space, posing challenges in adapting it effectively to latent diffusion frameworks. Another line of work is fine-tuning. In Miao_2024_CVPR , the authors provide a finetuning method using Reinforcement Learning to improve the diversity of generated samples using a diversity reward function.
@@
Recent works tackle this problem in the denoising phase. The CADS framework sadat2024cads shows that adding Gaussian noise to the conditioning signals during inference increases sample diversity. Particle Guidance (PG) corso2024particleguidance employs non-IID sampling from the joint distribution defined by a diffusion model combined with a potential function that maximizes pairwise dissimilarity across all samples, independent of semantic context. Similar to PG, ProCreate lu2024procreate uses DreamSim embeddings to find similar images via log energy and maximizes embedding-space distances. The concurrent method SPELL kirchhof2024sparse adds repellency terms during sampling to prevent samples in a batch from being too close. These methods maximize across all samples without considering the prompts. However, these methods are prompt-agnostic, while ours dynamically conditions diversity guidance on the input.
@@
The recent work askari2024improving introduces contextualized Vendi Score Guidance (c-VSG), enhancing generative diversity during the denoising process using Vendi Score friedman2023vendi . Their approach, however, requires the same prompts, restricting its applicability across diverse prompt scenarios. On the other hand, our method leverages the RKE Score jalali2023information to significantly enhance computational efficiency and sample complexity. Additionally, we propose prompt-aware guidance inspired by jalali2024conditional , enabling adaptive soft-clustering and effective conditioning on semantically distinct prompts. Moreover, unlike askari2024improving , which applies guidance based on latent encoded features of reference images, our strategy directly applies guidance within the diffusion model’s latent space, saving computational costs and enabling a prompt-aware diversity improvement.
@@ 3 Preliminaries
3.1 Kernel function and Vendi diversity scores
@@
We assume that the kernel function is normalized, i.e., $k(x,x)=1$ for every $x\in\mathcal{X}$. For a general (potentially unnormalized) kernel $k$, one can define its normalized counterpart as $\widetilde{k}(x,y)=k(x,y)/\sqrt{k(x,x)\,k(y,y)}$, which we apply whenever normalization is required. The Rényi Kernel Entropy (RKE) diversity score jalali2023information is defined and analyzed as the inverse Frobenius norm-squared of the trace-normalized kernel matrix:
$\mathrm{RKE}\bigl(x_{1},\ldots,x_{n}\bigr):=\exp\!\Bigl(H_{2}\bigl(\tfrac{1}{n}K_{X}\bigr)\Bigr)\,=\,\Bigl\|\tfrac{1}{n}K_{X}\Bigr\|_{F}^{-2}$ (2)
where $\|\cdot\|_{F}$ denotes the Frobenius norm. As theoretically shown in jalali2023information , the RKE score can be regarded as a mode count of a mixture distribution with multiple modes.
To extend this entropy-based diversity scores to conditional (i.e., prompt-aware) diversity measurement for prompt-guided generative models, jalali2024conditional propose the application of conditional kernel matrix entropy measures, resulting in the following definition of order-$2$ Conditional-RKE given the Hadamard product $\frac{1}{n}K_{X}\odot K_{T}$ of the output data $K_{X}$ and input prompt $K_{T}$ kernel matrices:
$\mathrm{Conditional}\text{-}\mathrm{RKE}\bigl(x_{1},\ldots,x_{n};t_{1},\ldots,t_{n}\bigr)\,:=\,\frac{\bigl\|K_{T}\bigr\|_{F}^{\,2}}{\bigl\|K_{X}\odot K_{T}\bigr\|_{F}^{\,2}},$ (3)
Note that here the Hadamard product of the prompt $(t_{1},\ldots t_{n})$ kernel matrix $K_{T}$ and output $(x_{1},\ldots x_{n})$ kernel matrix $K_{X}$ is defined as the elementwise product of the matrices, which is guaranteed to be a PSD matrix by the Schur product theorem.
@@ 3.2 Conditional Latent Diffusion Models
Latent Diffusion Models (LDMs) rombach2022high can achieve scalability and efficiency by performing the diffusion process within a compressed latent space, rather than directly in pixel space. A pre-trained autoencoder is used in LDMs, consisting of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. The encoder $\mathcal{E}$ maps high-dimensional images $\boldsymbol{x}\in\mathbb{R}^{H\times W\times C}$ to a lower-dimensional latent representation $\boldsymbol{z}_{0}=\mathcal{E}(\boldsymbol{x})\in\mathbb{R}^{h\times w\times c}$, while the decoder $\mathcal{D}$ reconstructs the images $\tilde{\boldsymbol{x}}=\mathcal{D}(\boldsymbol{z}_{0})$ from a denoised latent variable $\boldsymbol{z}_{0}$. The forward diffusion process applies in the latent space as $\boldsymbol{z}_{t}=\sqrt{\alpha_{t}}\boldsymbol{z}_{0}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}_{t}$, where $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathit{\boldsymbol{I}})$ and $\alpha_{t}\in$ are predefined noise schedule parameters. Then the conditional LDMs with conditions $\boldsymbol{y}$ are trained to predict the noise $\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z}_{t},t,\boldsymbol{y})$ at each time step $t$, also learn the score of $p_{t}(\boldsymbol{z_{t}}|\boldsymbol{y})$ song2019generative ; song2020score :
$\min_{\theta}\mathbb{E}_{\boldsymbol{z}_{t},\boldsymbol{\epsilon}_{t},t,\boldsymbol{y}}\left[\left\|\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z}_{t},t,\boldsymbol{y})-\boldsymbol{\epsilon}_{t}\right\|_{2}^{2}\right]=\min_{\theta}\mathbb{E}_{\boldsymbol{z}_{t},\boldsymbol{\epsilon}_{t},t,\boldsymbol{y}}\left[\left\|\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z}_{t},t,\boldsymbol{y})+\sqrt{1-\alpha_{t}}\nabla_{\boldsymbol{z}_{t}}\log p_{t}(\boldsymbol{z}_{t}|\boldsymbol{y})\right\|_{2}^{2}\right].$ (4)